# LightPAL: Lightweight Passage Retrieval for Open Domain Multi-Document Summarization

### v1

@@ 5.1 Settings
Detailed Settings. In the proposed method, we used the bge-large-en-v1.5 embedding model to obtain the top-100 most similar passages for each passage as candidates. For a fair comparison with PromptRank, we calculated the ContextScore for these passage pairs using Mistral-7B-v0.1, the same model employed by PromptRank. We then constructed a graph by creating edges between the top-5 pairs based on their ContextScore. The ContextScore computation, performed using A100 PCIe 80GB (8 GPUs), took 43 hours for QsumOD, 97 hours for Meeting, and 9 hours for Story datasets. In the Personalized PageRank algorithm, edge weights were uniformly set, and the algorithm was configured to jump uniformly to the top-20 of initial passages $D_{init}$ with a probability of $1-\alpha$, where $\alpha$ was set to 0.2.
Command-r Mixtral
Comp Div Rel Comp Div Rel
@@
PrR 66 67 60 55 60 57
PAL 56 55 55 49 49 57
QsumOD
PrR 51 58 53 53 57 48
PAL 53 57 44 54 54 42
@@
PrR 40 47 53 55 60 52
PAL 45 53 39 45 50 53
Table 2: Head-to-head win rates vs. a naive dense retriever for summary quality (Comprehensiveness, Diversity, and Relevance w.r.t reference summary) judged by GPT-4. <50% in red, >50% in blue. Results when the maximum number of retrieved passages are shown. PrR and PAL stand for PromptRank and LightPAL, respectively.
Language Models for Summary Generation. For the ODSUM task, where the number of passages required for generating high-quality summaries is unknown in advance, we used top-tier long-context language models from the RULER benchmark Hsieh et al. (2024) as summary generators to handle a wide range of context lengths: c4ai-command-r-v01  [^6] https://huggingface.co/CohereForAI/c4ai-command-r-v01 and mixtral-8x22B-Instruct-v0.1  [^7] https://huggingface.co/mistralai/Mixtral-8x22B-Instruct-v0.1. For Command-r, we evaluated from 20 to 200 passages for Meeting, from 10 to 100 passages for Story, and from 20 to 100 passages for QsumOD. Due to context limits, Mixtral was evaluated with up to 40 passages for Story and 60 for QsumOD. The prompts used for summary generation are provided in Appendix A.2.
Evaluation Aspects. We conducted a pairwise comparison using an LLM  [^8] We used OpenAI’s gpt-4-turbo-2024-04-09 (https://platform.openai.com/docs/models/gpt-4-turbo-and-gpt-4) as the evaluator. to investigate whether our proposed method and PromptRank generate higher-quality summaries than the naive approach for the same queries. The LLM greedily decoded and chose between two options (A or B) for the better summary. We evaluated the summaries along the following three dimensions: (1) Relevance: How much important content from the reference summaries is covered. Assessed whether the summary captures the key points and essential information from the reference summaries. (2) Diversity: How varied and insightful the summary is in providing different perspectives on the question. Assessed whether the summary introduces multiple viewpoints and facilitates a broader understanding of the topic. (3) Comprehensiveness: How well the summary covers all aspects of the question. Evaluated whether the summary provides a thorough representation of information that directly addresses the question. The prompts used for evaluation are provided in Appendix A.1.
CPU time GPU time
Data LightPAL LightPAL PromptRank
ST 0.001 s 0.014 s 46.506 s
QS 0.007 s 0.016 s 17.993 s
MT 0.017 s 0.017 s 12.342 s
Table 3: Average latency per query in context passage retrieval in seconds. The CPU used is an AMD EPYC 7302 16-Core Processor (one thread), and the GPU is an A100 PCIe 80GB (8 GPUs). MT, ST, and QS are abbreviations for Meeting, Story, and QsumOD, respectively.
Figure 2: Average win rates by graph construction method (GenProb: Language Model Generation Probability, EmbSim: Passage Embedding Similarity) when the summary generation model is fixed to Command-r. Top: Passage Retrieval w/ LightPAL. Bottom: Passage Retrieval w/ PromptRank.
@@ 5.2 Summary Quality Comparison
Table 2 presents the pair-wise comparison of summaries against the naive retriever. Story: The proposed and existing methods outperform the naive dense retriever, with PromptRank having a significant impact (up to 67% win rate for div). LightPAL shows smaller but consistent improvements, except for Mixtral’s Comp/Div, which are mostly tied. QsumOD: Both the proposed and existing methods improve Comprehensiveness and Diversity over the naive dense retriever, especially for Command-r. However, Relevance to reference summaries generally deteriorates, possibly due to the short reference summaries (see Table 1) lacking diverse content relative to the query, resulting in a trade-off between diversity and relevance. Meeting: In general, both the LightPAL and PromptRank result in worse performance compared to the naive dense retriever. This may be attributed to two factors. First, the queries in the Meeting dataset are artificially composed of multiple specific queries in QMSum dataset Zhong et al. (2021). The dense retriever might have already retrieved sufficiently query-relevant passages via such specific queries. Second, the Meeting dataset contains passages in a direct speech style, which may make it more challenging to capture the dependencies between the content. Utterances tend to be more verbose than third-person narrative text, and this verbosity may hinder the understanding of content dependencies. The impact of graph construction methods on performance will be explained in Section 5.4, and the retrieval performance will be discussed in Section 5.5.
In conclusion, PromptRank and LightPAL improve performance in QsumOD and Story datasets, particularly in Comprehensiveness and Diversity. The proposed method, while inferior to PromptRank, still improves performance favorably.
5.3 Latency Comparison in Passage Retrieval
Table 3 shows the average latency for retrieving context passages for a single query. On both CPU and GPU, LightPAL is approximately 1,000 times faster than PromptRank. Also, PromptRank’s latency exceeds 10 seconds, which is prohibitive for interactive RAG systems. The time complexity for computing Personalized PageRank scores is linear in the number of edges in the graph Yang et al. (2024), allowing for practical search time.
Story (top-$k$) QsumOD (top-$k$)
Method 10 20 30 40 50 20 40 60 80 100
Precision@$k$
Naive 29.62 17.54 13.00 10.38 8.67 14.57 9.81 7.50 6.04 5.13
PrR 48.85 32.00 23.49 18.27 14.97 13.35 9.72 8.74 7.61 6.62
PAL 43.15 25.67 17.92 14.29 11.92 13.78 9.31 7.43 6.13 5.14
Recall@$k$
Naive 36.61 42.28 46.81 49.41 51.27 37.05 47.48 52.84 56.40 58.89
PrR 56.52 70.86 77.03 79.88 82.01 33.42 43.36 49.60 52.93 55.27
PAL 50.54 58.45 60.66 64.61 67.39 34.98 44.96 52.95 56.87 59.27
Meeting (top-$k$)
Method 20 40 60 80 100
Precision@$k$
Naive 19.69 16.15 14.39 13.34 12.46
PrR 22.06 19.47 16.82 14.94 13.68
PAL 21.37 18.07 15.89 14.55 13.69
Recall@$k$
Naive 3.02 4.90 6.35 7.67 8.77
PrR 3.49 5.95 7.50 8.77 9.80
PAL 3.28 5.52 7.17 8.56 9.79
Table 4: Comparison of retrieval performance for each method with varying numbers of total passages. PrR and PAL stand for PromptRank and LightPAL, respectively.
5.4 Impact of Graph Construction Method
We analyze the impact of graphs constructed using LLM generation probabilities (GenProb) on summarization quality. As a baseline, we create a graph (EmbSim) that connects each passage to the top-5 passages with the highest cosine similarity based on passage embeddings by bge-large-en-v1.5. Both graphs have the same number of edges per passage.
Figure 2 shows the average win rate of both graphs against the naive retriever in terms of Comprehensiveness, Diversity, and Relevance. Regardless of whether PromptRank or LightPAL is used for context passage retrieval, the GenProb graph outperforms EmbSim on the Story dataset. This is likely because Story passages are sections of a cohesive narrative, and retrieving the entire story is necessary to answer queries. The language model captures the clear sequential relationship between sections of the same story. Furthermore, in PromptRank, the GenProb graph tends to outperform EmbSim on other datasets (Meeting, QsumOD), while this is not the case with LightPAL. These results suggest that the effectiveness of graph construction methods depends on (1) the structure of the passage set and (2) the graph traversal algorithm. At least in PromptRank, which performs iterative reasoning, linking passages using language model generation probabilities (GenProb) positively impacts retrieval.
5.5 Retrieval Performance
To confirm that the improvement in summary quality is due to enhanced retrieval performance, we evaluated Precision@$k$ and Recall@$k$ for the top-$k$ retrieved passages. Precision@$k$ measures the fraction of passages among the top $k$ retrieved results that are actually relevant to the query. Recall@$k$, on the other hand, measures the fraction of all relevant passages that are successfully retrieved within the top $k$ results. Better retrieval performance leads to retrieved passages being more aligned with the reference summary since it was created using query-relevant passages. Consequently, the generated summary from retrieved passages becomes more consistent with the reference summary, which will manifest as an increase in the Relevance metric.
Table 4 presents results when varying the number of top-$k$ passages. Story: LightPAL/PromptRank shows clear improvement over the naive approach, aligning with the improvement in Relevance seen in Section 5.2. QsumOD: LightPAL/PromptRank exhibits little variation from the naive method, which is consistent with the lack of improvement in Relevance. Meeting: LightPAL/PromptRank demonstrates a marginal performance improvement. However, Section 5.2 confirmed that summary quality, especially Relevance, does not improve from the naive approach. This result suggests that our settings to labeling relevant passages for the Meeting dataset may be flawed. Dividing the relevant meeting script into smaller passages and labeling all of them as relevant has likely led to marking many irrelevant passages as relevant, potentially resulting in an inaccurate evaluation of the retrieval performance.
@@ 6 Conclusion
In this paper, we proposed LightPAL, a lightweight passage retrieval method for Open-Domain Multi-Document Summarization. Our method pre-constructs a graph representing passage relationships using language models during indexing. At retrieval time, random walk via Personalized PageRank on this graph enables low-latency retrieval of diverse relevant information without runtime LM inference. Experiments on ODMDS and query-focused summarization benchmarks show that our proposed method outperforms baseline retrievers in summary quality while being significantly more efficient than an iterative retrieval and reasoning approach for Multi-hop QA.
@@ 7 Limitations
Lack of human evaluation. This paper lacks human evaluation of the generated summaries. Although LLM-based summary evaluations are known to align moderately with human judgments Liu et al. (2023), actual human evaluations are required to further validate the reliability of the results.
Evaluation on datasets other than ODMDS benchmarks. This research has only been evaluated on benchmarks for ODMDS and query-focused multi-document summarization. However, there are other tasks that require long answers based on queries for large document collection. A prominent example is long-form QA Fan et al. (2019), where the queries are complex and open-ended.
Variation of retrieval model (dense/sparse). Our proposed method can use any retrieval model for initial passage retrieval. This paper used a single SOTA dense retriever, i.e., bge-large-en-v1.5. However, the impact of using other pretrained dense retrievers or sparse retrievers like BM25 has not been evaluated.
Variation of LLMs for summary generation. This paper used representative LLMs capable of handling long contexts (c4ai-command-r-v01 and mixtral-8x22B-Instruct-v0.1) and evaluated the quality of summaries when inputting the maximum possible number of passages. However, this methodology increases the inference cost of LLMs. A more practical approach would be to limit the number of retrieved passages, i.e., reduce information recall and increase precision, for summary generation. In such scenarios, models with shorter context lengths may be sufficient for generating summaries. Therefore, a comprehensive evaluation using diverse language models as summarizers is required.
Variation of language model for PromptRank and other iterative approaches. PromptRank is not the only approach that iteratively performs reasoning and retrieval. While PromptRank was chosen as a representative approach for comparison, other methods using Chain-of-Thoughts Trivedi et al. (2023) or LLMs as graph traversal agent Wang et al. (2024) exist. Moreover, there is a trade-off between the capability (size) and inference speed of the language model used for reasoning. Optimizing this trade-off could lead to faster methodologies, but this evaluation has not been performed.

### v2

@@ 5.1 Settings
Detailed Settings of LightPAL. For graph construction of LightPAL, we used the bge-large-en-v1.5 embedding model to obtain the top-100 most similar passages for each passage as candidates. For a fair comparison with PromptRank, we calculated the ContextScore for these passage pairs using the same language models employed by PromptRank. We then constructed a graph by creating edges between the top-5 pairs based on their ContextScore. In the Personalized PageRank algorithm, edge weights were uniformly set, and the algorithm was configured to jump uniformly to the top-20 of initial passages $D_{init}$ with a probability of $1-\alpha$, where $\alpha$ was set to 0.2.
Number of Passages and Language Models for Summary Generation. For the ODSUM task, we speculate that the number of retrieved passages affects the final summary quality. Thus, we tested three scenarios: the retrieved passages are fewer than, approximately equal to, or more than the average number of relevant passages in each dataset (see Table 1 for detailed statistics). In this way, we experimented with multiple retrieval settings: 5, 10, and 20 passages for Story and RAGQA data, and 100, 200, and 300 ones for Meeting data. Settings with higher numbers require processing larger volumes of text. To handle this long-context setting, we employed top-tier long-context LLMs from the RULER benchmark Hsieh et al. (2024) as our summary generators: Meta-Llama-3.1-8B-Instruct [^5] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct and MegaBeam-Mistral-7B-512k [^6] https://huggingface.co/aws-prototyping/MegaBeam-Mistral-7B-512k. The prompts used for summary generation are provided in Appendix A.2.
Method Story Meeting RAG-QA
Pre@k Rec@k Pre@k Rec@k Pre@k Rec@k
bge-large-en-v1.5 31.15 36.27 10.34 12.79 64.25 71.49
+PromptRank (Qwen 0.5B) 31.51 37.36 12.22 11.16 59.77 67.71
+PromptRank (Qwen 1.5B) 34.96 40.77 10.94 12.70 59.99 68.06
+PromptRank (Qwen 3B) 36.37 42.24 11.48 12.52 59.89 68.24
+LightPAL (Qwen 0.5B) 37.16 42.73 11.37 14.26 62.78 70.58
+LightPAL (Qwen 1.5B) 38.50 44.24 11.41 14.32 63.14 70.73
+LightPAL (Qwen 3B) 39.55 45.30 11.47 14.38 63.43 70.84
BM25 43.85 49.76 11.04 13.36 50.23 54.91
+PromptRank (Qwen 0.5B) 40.04 46.92 12.27 12.85 53.34 59.67
+PromptRank (Qwen 1.5B) 42.63 49.31 11.66 13.43 54.00 60.63
+PromptRank (Qwen 3B) 43.25 50.12 12.44 13.54 53.75 60.37
+LightPAL (Qwen 0.5B) 48.75 55.05 12.30 15.17 54.58 60.90
+LightPAL (Qwen 1.5B) 49.72 55.94 12.18 15.13 54.95 61.42
+LightPAL (Qwen 3B) 50.22 56.47 12.36 15.30 55.66 61.88
Table 2: Retrieval performance comparison among PromptRank and LightPAL with dense (bge-large-en-v1.5) and sparse (BM25) retrievers. Metrics include Precision@k and Recall@k (abbreviated as Pre@k and Rec@k in the table). Within each group, the highest score is shown in bold and underlined; the second-highest score is in bold. Scores are averaged across different numbers of retrieved passages.
5.2 Retrieval Performance Evaluation
Evaluation Metric. To first confirm that our proposed method enhances retrieval performance, we evaluated Precision@$k$ and Recall@$k$ for the top-$k$ retrieved passages. Precision@$k$ measures the fraction of passages among the top-$k$ retrieved results relevant to the query, and Recall@$k$ does the fraction of all relevant passages successfully retrieved within the top-$k$ results.
Effectiveness. Both LightPAL and PromptRank utilize LLMs, and we hypothesized that the retrieval performance might vary depending on the capacity of these models. To investigate this, we conducted experiments using models of different sizes from the same series: Qwen2.5-0.5B, Qwen2.5-1.5B, and Qwen2.5-3B [^7] https://github.com/QwenLM/Qwen2.5. Table 2 shows LightPAL’s retrieval performance using dense or sparse retrievers as base models. LightPAL improves performance for most dataset and base model combinations, except for RAG-QA and bge-large-en-v1.5. Performance enhancement is most significant for Story data, improving up to 10%. LightPAL performs comparably or better than PromptRank, outperforming it by 3-6% on Story/RAG-QA data. The experimental results also demonstrated the impact of LLM size on retrieval performance in LightPAL and PromptRank: using larger LLM models gradually improved retrieval performance, especially for Recall@$k$ metrics. However, performance gains from larger models are relatively small compared to improvements over baseline models. This suggests that even models smaller than 0.5B may still offer substantial benefits.
Search Latency. Table 3 presents the average latency for retrieving initial passages and context passages for a single query. The results demonstrate that LightPAL is approximately 1,000 times faster than PromptRank. Moreover, PromptRank’s latency exceeds 10 seconds, which is prohibitively slow for interactive RAG systems. In contrast, LightPAL’s latency is only about 10 to 20 times higher than the initial search’s. This suggests that LightPAL is unlikely to impair the real-time performance of RAG systems. The time complexity for computing Personalized PageRank scores is linear in the number of edges in the graph Yang et al. (2024), which allows for practical search times.
Graph Construction Time. Table 4 shows the time required for graph construction in LightPAL. These results demonstrate that graph construction can be achieved through lightweight processing. For instance, using Qwen2.5-0.5B on the Meeting dataset (approximately 10k passages), the process completes in about 4 hours on 8 NVIDIA L40S GPUs. Our results show that LightPAL improves retrieval performance while maintaining low search latency through offline graph construction.
Init. PR [ms]
data bge-large-en-v1.5 BM25
Story 2.23 1.59
Meeting 5.24 32.31
RAG-QA 2.50 2.89
Cont. PR [ms]
PromptRank LightPAL
Story $7.57\times 10^{4}$ 6.02
Meeting $3.09\times 10^{4}$ 113.31
RAG-QA $1.96\times 10^{4}$ 23.09
Table 3: Average search latency for Initial Passage Retrieval (Init. PR) and Context Passage Retrieval (Cont. PR). For Init. PR, latency includes both embedding and search time. Query embedding for bge-large-en-v1.5 was performed on NVIDIA L40 GPU. Search operations for bge-large-en-v1.5, BM25, and LightPAL were conducted on a single thread of AMD EPYC 9554 processor. For PromptRank, LLM inference (using Qwen2.5-3B model) was executed on a single NVIDIA L40S GPU.
Story Meeting RAG-QA
Qwen2.5-0.5B 0h 22m 3h 43m 0h 57m
Qwen2.5-1.5B 0h 44m 7h 40m 1h 58m
Qwen2.5-3B 1h 11m 12h 30m 3h 13m
Table 4: Computational time to calculate ContextScore for each language model in LightPAL. The preceding and following passages were truncated to a combined length of 1024 tokens. Inference was performed using data parallelism across 8 NVIDIA L40S GPUs.
Method Rel Comp Div ROUGE
@@ 5.2 Retrieval Performance Evaluation
bge-large-en-v1.5 2.34 3.46 2.14 10.86
+PromptRank 2.47 3.47 2.11 11.20
+LightPAL 2.44 3.50 2.16 10.99
bm25 2.56 3.58 2.23 11.54
+PromptRank 2.59 3.57 2.17 11.56
+LightPAL 2.59 3.60 2.22 11.53
@@
bge-large-en-v1.5 1.79 3.33 2.69 11.95
+PromptRank 1.82 3.36 2.73 12.21
+LightPAL 1.78 3.34 2.71 11.87
bm25 1.79 3.20 2.64 11.93
+PromptRank 1.81 3.29 2.66 12.30
+LightPAL 1.77 3.23 2.66 11.95
RAG-QA
bge-large-en-v1.5 2.70 3.81 3.38 19.16
+PromptRank 2.63 3.75 3.34 19.03
+LightPAL 2.71 3.80 3.36 19.17
bm25 2.55 3.62 3.10 18.06
+PromptRank 2.63 3.70 3.21 18.72
+LightPAL 2.64 3.66 3.19 18.56
Table 5: Comparison of summary quality across different datasets and base retrievers. Scores that improved from the base retriever are displayed in blue, while others are in red. Summary quality metrics (Relevance, Comprehensiveness, and Diversity) are evaluated using G-Eval Liu et al. (2023), which assigns weights to scores 1-5 based on GPT-4’s generation probabilities. The displayed scores are averaged across various numbers of retrieved passages and language models used for summary generation.
Dataset Rel Comp Div
$\rho$ $\tau$ $\rho$ $\tau$ $\rho$ $\tau$
Story .47 .36 .49 .40 .50 .38
RAG-QA .36 .27 .48 .39 .18 .13
Table 6: Summary-level correlation coefficients (Spearman’s $\rho$ and Kendall’s $\tau$) between human judgments and GPT-4 judgments for different datasets and aspects.
@@ 5.3 Summary Quality Comparison
Evaluation Metric. To evaluate the generated summaries, we calculated the average of ROUGE-1, -2, and -L scores for relevance to reference summaries. We also used G-Eval Liu et al. (2023), which weights scores 1-5 based on GPT-4’s generation probabilities, to assess three aspects: (1) Relevance: How much important content from the reference summaries is covered. Assessed whether the summary captures the key points and essential information from the reference summaries. (2) Diversity: How varied and insightful the summary is in providing different perspectives on the question. Assessed whether the summary introduces multiple viewpoints and facilitates a broader understanding of the topic. (3) Comprehensiveness: How well the summary covers all aspects of the question. Evaluated whether the summary provides a thorough representation of information that directly addresses the question. The prompts used for evaluation are provided in Appendix A.1.
Results. The comparison results of Summary Quality are presented in Table 5. These results demonstrate that LightPAL improves summarization performance in cases where retrieval performance was notably enhanced. While slightly inferior to PromptRank, LightPAL still shows improvements in overall summary quality. For the Story dataset, LightPAL achieves noticeable enhancements, particularly with bge-large-en-v1.5 base retriever. Both Relevance and ROUGE scores increase by approximately 0.1 points, aligning with the substantial improvement in retrieval performance. In contrast, the Meeting dataset shows only marginal improvements with LightPAL, consistent with minimal retrieval performance enhancement. For the RAG-QA dataset, using bm25 as the base retriever improves summarization performance: Relevance and Diversity increase by about 0.1 points, and ROUGE by 0.5 points. However, with the bge-large-en-v1.5 base retriever, performance remains largely unchanged or slightly decreases, aligning with the lack of retrieval performance improvement.
Human-GPT-4 Evaluation Correlation To verify the soundness of the GPT-4-based evaluation in the aforementioned results, we conducted human annotation on a subset of the dataset using identical criteria and calculated the correlation with GPT-4-based scores. We randomly sampled 20 queries each from the Story and RAG-QA datasets. We chose BM25 as the baseline search model and selected the number of passages that yielded the most significant performance boost for LightPAL. For each query, summaries generated by MegaBeam-Mistral-7B using BM25, BM25+LightPAL, and BM25+PromptRank were randomly shuffled and annotated by three natural language processing researchers. Two annotators each evaluated 10 different queries per dataset, while the third annotator covered all queries, resulting in two annotations per query. We then calculated the correlation between the average of these scores and the GPT-4-based scores. The annotation instructions and interface are detailed in Appendix B. Table 6 presents the correlation coefficients between human judgment and GPT-4 evaluation. Except for the RAG-QA and Div combination, we observe Spearman’s $\rho$ values ranging from approximately 0.3 to 0.5 across the other conditions. These moderate correlations suggest that GPT-4-based evaluations demonstrate a reasonable alignment with human judgments and its soundness.
Figure 2: Summary quality plot for samples with improved retrieval by LightPAL (Improved Set) versus those without (Non-improved Set). Each group represents a combination of dataset and base retriever, with the percentage of samples showing improved retrieval performance noted below.
5.4 Detailed Analysis
Does better retrieval performance really mean better summary quality? To improve the quality of summaries in ODMDS task, we proposed a method to enhance retrieval performance. But does better retrieval truly lead to higher quality summaries? To answer this question, we divided samples from Story and RAG-QA datasets into an Improved Set (where LightPAL improved Recall@$k$) and a Non-Improved Set. Figure 2 shows the results of our analysis. The Improved Set (shown in blue bars) demonstrates enhanced summary quality through LightPAL, particularly in terms of Relevance. Conversely, the Non-Improved Set shows either no change or a decline in performance. Notably, up to 62% of samples in the RAG-QA dataset with BM25 retriever showed improved retrieval performance, indicating that a significant number of samples require iterative retrieval method like LightPAL and PromptRank.
Effective Cases in LightPAL Retrieval Performance. To analyze when LightPAL succeeds or fails across various queries, documents, and base retrieval models, we present two cases in Figure 3. LightPAL enhances retrieval by 1) covering passages without distinctive query keywords and 2) selecting topically similar passages when multiple passages share common keywords. The latter is particularly beneficial compared to sparse retrievers like BM25. However, LightPAL may potentially degrade performance when distinctive keywords are present in both the target passage and the query, or when the system can accurately grasp the topic without assistance. For examples of generated summaries, please refer to Appendix C.
CASE 1: passages do not contain named entities in the query Query: What is the dynamic between Captain Midas and Mister Spinelli in the story of CAPTAIN MIDAS? Naive (bge-large-en-v1.5): ✗: The captain’s voice. Calm, brief. It sent a tremor… (a section of different story) LightPAL: ✓: The first thing about the derelict…
CASE 2: misunderstanding of the topic Query: How to see how much time you spend on a game on xbox? Naive (bm25): ✗: you have to get Nintendo’s permission, buy from Nintendo’s factories, …. If you want onto the XBox, sure there’s XBLA… LightPAL: ✓: go to the friends tab and the choose…
Figure 3: Case study of retrieved passages using LightPAL and a naive base retriever.
@@ 6 Conclusion
In this paper, we proposed LightPAL, a lightweight passage retrieval method for Open-Domain Multi-Document Summarization. Our method constructs a graph representing passage relationships using language models during indexing. At retrieval time, random walk via Personalized PageRank on this graph enables low-latency retrieval of diverse relevant information without runtime LLM inference. Experiments on ODMDS and long-form QA benchmarks demonstrate that LightPAL outperforms naive sparse/dense retrievers in both retrieval and summarization metrics, while achieving higher efficiency compared to an iterative retrieval and reasoning approach for multi-hop QA.
@@ 7 Limitations
Variation of LLMs for summary generation. This paper used representative LLMs capable of handling long contexts (Meta-Llama-3.1-8B-Instruct and MegaBeam-Mistral-7B-512) for summary generation. A comprehensive evaluation using diverse language models as summarizers would be preferable to further validate the effectiveness of LightPAL and identify potential limitations.
Limited comparison with other iterative reasoning and retrieval approaches. PromptRank Khalifa et al. (2023) is not the only approach that iteratively performs reasoning and retrieval. While PromptRank was chosen as a representative approach for comparison, other methods exist, such as those using Chain-of-Thoughts Trivedi et al. (2023) or LLMs as graph traversal agents Wang et al. (2024b). A more comprehensive comparison with these alternative approaches could provide additional insights into the strengths and limitations of our proposed method.
Selective application of LightPAL. LightPAL does not universally enhance retrieval performance across all samples. Additionally, its effectiveness varies depending on the dataset and base retriever used. To maximize the utility of our proposed method, developing criteria for when to apply LightPAL would be beneficial. For example, if the similarity between the query and retrieved passages exceeds a predetermined threshold, we might conclude that the base retriever is adequate without applying LightPAL.
Acknowledgments
We thank Takuya Tamura (NEC Corporation) for the discussion of this research.